OpenStudy (anonymous):

PLEASE HELP! MEDAL REWARDED!! solve for y in the first equation. x − 3y = −6 2x − 7y = 10

OpenStudy (anonymous):

hello there

OpenStudy (anonymous):

Ok here you go x + 6 = 3y x - 3y + 6= 0 y = x/3 + 2

OpenStudy (anonymous):

Is that what your looking for?

OpenStudy (anonymous):

so the answer is 1/3x+2 ?

OpenStudy (anonymous):

Yes y = x/3 + 2